OpenStudy (anonymous):

Find the slope of the line connecting the following points. (6, -1) (12, -3)

OpenStudy (anonymous):

Calculate \(\Delta x\) and \(\Delta y\): \[\Delta x = 6-12 = -6\\ \Delta y = -1 - (-3) = 2\\ \mbox{slope is } \frac{\Delta y}{\Delta x}=2/-6=-1/3 \]

OpenStudy (anonymous):

thanks